# Starfish lemma via birational quasi-isomorphisms

### v2

@@ 5.3 Starfish lemma in factorial domains
If $x_{i}\in\mathbf{x}_{0}(\kappa)$, then $\mathcal{Q}(x_{i})$ is an invertible element of $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$. Since $\tilde{\mathbf{x}}_{0}(\kappa)$ consists of prime elements, if $f$ is a prime factor of $x_{i}$, then $\mathcal{Q}(f)$ is a Laurent monomial in $\tilde{\mathbf{x}}_{0}$ with a coefficient from $\mathbb{C}^{\times}$; hence, $f$ is a Laurent monomial in $\mathbf{x}_{0}(\kappa)$. Applying $\mathcal{Q}^{\prime}$ to $x_{i}$ and $f$, we see that $\mathcal{Q}^{\prime}(x_{i})$ is prime in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$, and therefore, $\mathcal{Q}^{\prime}(f)$ is an associate of $\hat{x}_{\kappa^{\prime}(i)}$ in $\hat{R}[\hat{\mathbf{x}}(\kappa^{\prime})^{\pm 1}]$; hence, $f$ is an associate of $x_{i}$ in $R$, and thus $x_{i}$ is prime. In a similar way, one proves that elements of $\mathbf{x}_{0}(\kappa^{\prime})$ are prime. ∎
Corollary 5.7.
In the setup of Proposition 5.6, if all the initial variables in $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are irreducible, then all the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime.
@@
By Proposition 5.6, the variables in $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are irreducible, and as in the proof of Proposition 4.15, the initial variables in $\hat{\mathbf{x}}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ are pairwise coprime in the respective rings. It follows from Proposition 5.5 that all the variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime. ∎
Proposition 5.8.
Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric types in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $\Sigma_{0}:=(\mathbf{x}_{0},B_{0})$, $\tilde{\Sigma}:=(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $\hat{\Sigma}_{0}:=(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
@@
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
@@
    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. If in addition,
  1. 4)
    The initial exchange matrices $B_{0}$, $\tilde{B}_{0}$ and $\hat{B}_{0}$ are of full rank;
  2. 5)
    For the star-shaped nerves $\tilde{\mathcal{N}}$ and $\hat{\mathcal{N}}$ centered at $\tilde{\Sigma}_{0}$ and $\hat{\Sigma}_{0}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{N}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{N}})\supseteq\hat{R}$.
Then $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})=\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})=\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})=R$.
@@
It follows from Corollary 5.7 that the initial variables in $\mathbf{x}_{0}$ are irreducible and pairwise coprime. The rest follows from Proposition 5.4. ∎
Remark 5.9.
As we mentioned in Remark 4.16, one can produce other versions of Proposition 5.8 via relaxing the conditions on the ground field and the groups of invertible elements of the rings.
@@ Appendix A Example in $\mathbb{C}[\operatorname{SL}_{3}]$
In this appendix, we illustrate our theory in the case of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ compatible with Poisson brackets from the Belavin-Drinfeld888 Given a set of simple roots $\Delta$, a Belavin-Drinfeld triple is a triple $\mathbf{\Gamma}:=(\Gamma_{1},\Gamma_{2},\gamma)$ where $\Gamma_{1},\Gamma_{2}\subset\Delta$ and $\gamma:\Gamma_{1}\rightarrow\Gamma_{2}$ is a nilpotent isometry. Nilpotency means that for every $\alpha\in\Gamma_{1}$ there is $k\geq 1$ such that $\gamma^{k}(\alpha)\notin\Gamma_{1}$. Belavin-Drinfeld triples parameterize the moduli space of factorizable quasi-triangular Poisson brackets on a connected simple complex algebraic group $G$ [1, 4]. A more general class of Poisson homogeneous brackets is parameterized by pairs of Belavin-Drinfeld triples $(\mathbf{\Gamma}^{r},\mathbf{\Gamma}^{c})$ [10]. This general context is not significant for our purpose, which is to provide explicit examples of birationally quasi-isomorphic cluster structures. class. A combinatorial recipe for constructing such cluster structures was first given in [10]. A general construction of quasi-isomorphisms for cluster structures in $\mathbb{C}[\operatorname{SL}_{n}]$ (in the context of Belavin-Drinfeld brackets) is available in [11].
@@
(a) Case of $\bar{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
(b) Case of $\tilde{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}}_{\text{st}})$.
(c) Case of $\hat{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{st}},\operatorname{\mathbf{\Gamma}})$.
@@
Figure 1: The initial quivers for $\bar{\mathcal{C}}$, $\tilde{\mathcal{C}}$, $\hat{\mathcal{C}}$ and $\mathcal{C}$; here, $\operatorname{\mathbf{\Gamma}}=(\{2\},\{1\},2\mapsto 1)$.

### v3

@@ 6.1 Transferring cluster and algebraic properties
As in the proof of Proposition 4.5, since the variables are algebraically independent over $\mathbb{C}$, we see that the marked variables in $\tilde{R}$ and $\hat{R}$ are pairwise coprime. By Proposition 5.6, it is enough to show that each $x_{i}\in\mathbf{x}_{0}(\kappa)$ is coprime with each $x_{j}\in\mathbf{x}_{0}(\kappa^{\prime})$. Indeed, assume that $f\in R$ is such that $f\ |\ x_{i}$ and $f\ |\ x_{j}$. Then $\mathcal{Q}(f)$ is invertible in $\tilde{R}[\tilde{\mathbf{x}}_{0}(\kappa)^{\pm 1}]$, and therefore, $f$ is a Laurent monomial in $\mathbf{x}_{0}(\kappa)$ with a coefficient from $\mathbb{C}^{\times}$. Applying $\mathcal{Q}^{\prime}$, we see that $f$ is also invertible in $\hat{R}[\mathbf{x}_{0}(\kappa^{\prime})^{\pm 1}]$, and since the marked variables in $\hat{R}$ are prime and pairwise coprime, we conclude that $f\in\mathbb{C}^{\times}$. Therefore, $x_{i}$ is coprime with $x_{j}$, and by Proposition 5.6, we conclude that all the marked variables are prime in $R$. ∎
The next result corresponds to Proposition 4.6.
Proposition.
Let $R$, $\tilde{R}$ and $\hat{R}$ be factorial Noetherian $\mathbb{C}$-algebras with no zero divisors and such that $R^{\times}=\tilde{R}^{\times}=\hat{R}^{\times}=\mathbb{C}^{\times}$, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended clusters $\mathbf{x}_{0}$, $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$. Assume the following:
  1. 1)
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
  2. 2)
    The initial extended clusters $\tilde{\mathbf{x}}_{0}$ and $\hat{\mathbf{x}}_{0}$ are i-admissible;
  3. 3)
    The initial extended cluster $\mathbf{x}_{0}$ is regular.
Then $\mathbf{x}_{0}$ is i-admissible, as well as $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$, $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$.
@@
It follows from Condition 2 and the Starfish lemma that $\bar{\mathcal{A}}_{\mathbb{C}}(\hat{\mathcal{C}})\subseteq\hat{R}$ and $\bar{\mathcal{A}}_{\mathbb{C}}(\tilde{\mathcal{C}})\subseteq\tilde{R}$. By Lemma 6.3, the marked variables in $R$ are irreducible and pairwise coprime. By Proposition 5.4, each mutation $x^{\prime}$ of a marked variable $x$ is regular and coprime with $x$; hence, $\mathbf{x}_{0}(\kappa,\kappa^{\prime})$ is i-admissible. It follows from Proposition 4.3 that $\mathbf{x}_{0}$ is admissible and from Proposition 4.5 that $\mathbf{x}_{0}$ is i-admissible and $\bar{\mathcal{A}}_{\mathbb{C}}(\mathcal{C})\subseteq R$. ∎
6.4 Proof of Proposition 4.11
We recall the statement of Proposition 4.11 with the full setup.
Proposition.
Let $R$, $\tilde{R}$ and $\hat{R}$ be normal Noetherian domains, $(\mathcal{C},R)$, $(\tilde{\mathcal{C}},\tilde{R})$ and $(\hat{\mathcal{C}},\hat{R})$ be cluster structures of geometric type in $R$, $\tilde{R}$ and $\hat{R}$, with fixed initial extended seeds $(\mathbf{x}_{0},B_{0})$, $(\tilde{\mathbf{x}}_{0},\tilde{B}_{0})$ and $(\hat{\mathbf{x}}_{0},\hat{B}_{0})$. Assume the following:
@@
    There are markers $\kappa$ and $\kappa^{\prime}$ that make $(\mathcal{C},\tilde{\mathcal{C}},\kappa)$ and $(\mathcal{C},\hat{\mathcal{C}},\kappa^{\prime})$ related triples, as well as a pair of complementary birational quasi-isomorphisms $\mathcal{Q}:(\mathcal{C},R)\rightarrow(\tilde{\mathcal{C}},\tilde{R})$ and $\mathcal{Q}^{\prime}:(\mathcal{C},R)\rightarrow(\hat{\mathcal{C}},\hat{R})$;
@@
    The matrices $\tilde{B}_{0}$, $\hat{B}_{0}$ and $B_{0}$ are of full rank;
  3. 3)
    The inclusions $\bar{\mathcal{A}}(\tilde{\mathcal{C}})\supseteq\tilde{R}$ and $\bar{\mathcal{A}}(\hat{\mathcal{C}})\supseteq\hat{R}$ hold.
Then $\bar{\mathcal{A}}(\mathcal{C})\supseteq R$.
@@ Appendix A Example in $\mathbb{C}[\operatorname{SL}_{3}]$
In this appendix, we illustrate our theory in the case of cluster structures in $\mathbb{C}[\operatorname{SL}_{3}]$ compatible with Poisson brackets from the Belavin-Drinfeld777 Given a set of simple roots $\Delta$, a Belavin-Drinfeld triple is a triple $\mathbf{\Gamma}:=(\Gamma_{1},\Gamma_{2},\gamma)$ where $\Gamma_{1},\Gamma_{2}\subset\Delta$ and $\gamma:\Gamma_{1}\rightarrow\Gamma_{2}$ is a nilpotent isometry. Nilpotency means that for every $\alpha\in\Gamma_{1}$ there is $k\geq 1$ such that $\gamma^{k}(\alpha)\notin\Gamma_{1}$. Belavin-Drinfeld triples parameterize the moduli space of factorizable quasi-triangular Poisson brackets on a connected simple complex algebraic group $G$ [1, 4]. A more general class of Poisson homogeneous brackets is parameterized by pairs of Belavin-Drinfeld triples $(\mathbf{\Gamma}^{r},\mathbf{\Gamma}^{c})$ [10]. This general context is not significant for our purpose, which is to provide explicit examples of birationally quasi-isomorphic cluster structures. class. A combinatorial recipe for constructing such cluster structures was first given in [10]. A general construction of quasi-isomorphisms for cluster structures in $\mathbb{C}[\operatorname{SL}_{n}]$ (in the context of Belavin-Drinfeld brackets) is available in [11].
@@
(a) Case of $\bar{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{std}},\operatorname{\mathbf{\Gamma}}_{\text{std}})$.
(b) Case of $\tilde{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}},\operatorname{\mathbf{\Gamma}}_{\text{std}})$.
(c) Case of $\hat{\mathcal{C}}$, $(\operatorname{\mathbf{\Gamma}}^{r},\operatorname{\mathbf{\Gamma}}^{c})=(\operatorname{\mathbf{\Gamma}}_{\text{std}},\operatorname{\mathbf{\Gamma}})$.
@@
Figure 1: Initial quivers for $\bar{\mathcal{C}}$, $\tilde{\mathcal{C}}$, $\hat{\mathcal{C}}$ and $\mathcal{C}$; here, $\operatorname{\mathbf{\Gamma}}=(\{2\},\{1\},2\mapsto 1)$ and $\operatorname{\mathbf{\Gamma}}_{\text{std}}=(\emptyset,\emptyset,\emptyset\rightarrow\emptyset)$.